OpenStudy (nathanjhw):

If 12 ft2 of material is available to make a box with a square base and an open top, find the largest possible volume of the box.

OpenStudy (nathanjhw):

@jim_thompson5910

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

lets call the base \(x\) and the height \(h\) so that the volume is \(V=x^2h\)

OpenStudy (misty1212):

then the total area is \[x^2+4xh\] which you know is 12 so set \[x^2+4xh=12\] solve for \(h\)in terms of \(x\) and put it back in the the equation of the volume

OpenStudy (misty1212):

you good from there?

OpenStudy (nathanjhw):

I'm still confused.

OpenStudy (nathanjhw):

@misty1212

OpenStudy (nathanjhw):

I think the answer is 2ft^3 is that correct?

OpenStudy (nathanjhw):

@misty1212

OpenStudy (misty1212):

i have no idea

OpenStudy (misty1212):

seem unlikely it is a whole number , but it could be

OpenStudy (nathanjhw):

The possibilities are 2 ft3 4 ft3 5 ft3 8.5 ft3 9 ft3

OpenStudy (misty1212):

\[x^2+4xh=12\\ 4xh=12-x^2\\ h=\frac{12-x^2}{4x}\] so \[V(x)=x^2\times\frac{12-x^2}{4x}\]

OpenStudy (misty1212):

clean that up, take the derivative etc

OpenStudy (nathanjhw):

When I solved for h and x I got integer solutions h= -1 , x = -2 ; h= -1, x= 6; h=1 , x=-6; h=1 , x =2

OpenStudy (nathanjhw):

I plugged it back in and got 2 and 6 and assumed 2 was the answer because it was one of the choices.

OpenStudy (nathanjhw):

@misty1212